OpenStudy (anonymous):

let y>0 and n a positive integer. Prove that (1+y)^n >= 1+ ny

OpenStudy (freckles):

Hey...@mhannahw are you here... Like I seen this post like 4 times now and I have responded to all three + this one now.

OpenStudy (anonymous):

Oh yeah sorry I'll look at this one I got confused

OpenStudy (freckles):

So the first positive integer is 1.

OpenStudy (freckles):

But the way the method we are going to prove this by is induction.

OpenStudy (freckles):

Like that is the method I'm familiar if anyways.

OpenStudy (freckles):

The base case is the first case (the case we first need to prove it for) Since we are trying to prove this for all positive integers That means we will need to start with n=1 Does the inequality hold for n=1?

OpenStudy (freckles):

\[(1+y)^1 \ge 1+(1)y \]

OpenStudy (freckles):

If this inequality is true then we move on to the next step of the proof.. The part that is actually kinda tough to really tough depending on what kinda algebra/some other matematical magic is involved

OpenStudy (freckles):

You suppose \[ 1+ky \le (1+y)^k \text{ for some integer } k>0 \\ \text{ And you want to show } 1+(k+1)y \le (1+y)^{k+1}\]

OpenStudy (freckles):

So starting with the 1+(k+1)y thing ...we want to show this is less than or equal to that other thing on the other side of the inequality and we want to some how do it by using what we have which is \[1+ky \le (1+y)^k\]

OpenStudy (freckles):

So given that you are starting with 1+(k+1)y ...can you somehow using the thing that we assumed was true...

OpenStudy (anonymous):

wait how did you get 1+(k+1)y≤(1+y)^k+1 (the second line when you were supposing some positive integer k)

OpenStudy (freckles):

Mathematical Induction: 1) Show for base case 2) Assume the statement holds for n=k, then show the statement holds for n=k+1

OpenStudy (anonymous):

Yeah but where did the 1+... and on the other side, the power k+1 come from in that line?

OpenStudy (freckles):

You are trying to prove... \[1+ny \le (1+y)^n \text{ right? } \]

OpenStudy (freckles):

I replaced the n's with (k+1)'s This inequality with (k+1)'s is the inequality we are trying to prove using the inequality with the k's.

OpenStudy (anonymous):

Okay, I see that. Can you explain why you replaced the n's with (k+1)'s?

OpenStudy (freckles):

Well {1,2,3,4,5,6,...,k,k+1} The way I always thought of it is... if we can see if the first case holds ... and then assume it holds for that k and show it holds for also k+1 then it holds for all the positive integers like this will give us all the positive integers like k could be 6 and k+1 could be 7 k could be 100 and k+1 could be 101 So what you are doing is showing it always holds for the next number in the set

OpenStudy (freckles):

Would you like me to go through the process with another inequality/equality...

OpenStudy (freckles):

Like I haven't proved your inequality here...I was trying to give you what you need to do it. But it seems like mathematical induction might actually be new to you.

OpenStudy (anonymous):

Precisely that! It is new to me. How would you show with a few cases that it works for every case?

OpenStudy (freckles):

the k (k>1) can be any number in the set with k+1 always following afterwards We are technically showing it for all positive integers

OpenStudy (freckles):

We aren't just pluggin in 6 and 7 or 91 and 100

OpenStudy (freckles):

but we are choosing k and k+1

OpenStudy (freckles):

which can represent any consecutive positive integer

OpenStudy (freckles):

integers*

OpenStudy (freckles):

\[1+2+3+\cdots +n=\frac{n(n+1)}{2} \text{ for all integer } n >0 \\ \text{ BASE CASE } \\ \text{ For } n=1: 1=\frac{1(1+1)}{2} \text{ holds ! } \\ \text{ ALL OTHER CASES (the induction case) } \\ \text{ Assume for some integer } k>0 \text{ we have } 1+2+3+\cdots +k=\frac{k(k+1)}{2} \\ \text{ Now we want to show } 1+2+3+\cdots +(k+1)=\frac{(k+1)(k+2)}{2} \\ \text{ So that task begins here } 1+2+3+\cdots+k+(k+1) \\ \\ = \frac{k(k+1)}{2}+k+1 \\ =\frac{k(k+1)}{2} +\frac{2(k+1)}{2} \\ =\frac{k(k+1)+2(k+1)}{2} \\ =\frac{(k+1)(k+2)}{2} \\ \text{ Therefore for all integer } n >0 \text{ we have }1+2+3+\cdots +n=\frac{n(n+1)}{2} \]\\

OpenStudy (freckles):

This above was an example of the induction proof.

OpenStudy (anonymous):

Right! okay I can follow that

OpenStudy (freckles):

The line after 1+2+3+...+k+(k+1) is where I used my induction hypothesis

OpenStudy (freckles):

aka the part we assumed was true

OpenStudy (freckles):

So in yours we had \[\text{ Induction hypothesis was: } 1+ky \le (1+y)^k \\ \text{ and we want to show } 1+(k+1)y \le (1+y)^{k+1}\]

OpenStudy (freckles):

We want to somehow reword the 1+(k+1)y so we can use what 1+ky is less than

OpenStudy (freckles):

Are you doing okay? Do you need a hint to start off?

OpenStudy (freckles):

I sleep now. Good luck.

OpenStudy (anonymous):

I think so, I'm finishing up! You helped much more than I could do on my own!

OpenStudy (anonymous):

Use the binomial Theorem to show that\[ (1+y)^n= 1+ ny +\text {something positive}\\ (y+1)^2=1+2y + y^2\\ (1+y)^3=1 + 3 y + 3 y^2 + y^3\\ (1+y)^4=1 + 4 y + 6 y^2 + 4 y^3 + y^4 \] and you are done